OpenStudy (anonymous):

simplify this expression: cscx/1+cscx - cscx/1-cscx

OpenStudy (anonymous):

\[\frac{\csc(x)}{1+\csc(x)}-\frac{\csc(x)}{1-\csc(x)}\] Is that what you mean?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{\csc(x)}{1+\csc(x)} - \frac{\csc(x)}{1-\csc(x)}\] \[\frac{\csc(x)[1-\csc(x)]}{[1+\csc(x)][1-\csc(x)]} - \frac{\csc(x)[1+\csc(x)}{[1-\csc(x)][1+\csc(x)]}\] \[\frac{\csc(x)-\csc ^{2}(x)-\csc(x)-\csc ^{2}(x)}{[1+\csc(x)][1-\csc(x)]}\] \[\frac{-2\csc ^{2}(x)}{[1-\csc(x)][1+\csc(x)}\] \[\frac{-2\csc ^{2}(x)}{1-\csc ^{2}(x)}\] \[\frac{-2\csc ^{2}(x)}{-\cot ^{2}(x)}\] \[\frac{\frac{2}{\sin ^{2}(x)}}{\frac{\cos ^{2}(x)}{\sin ^{2}(x)}}\] \[\frac{2}{\cos ^{2}(x)}\] \[2\sec ^{2}(x)\]

OpenStudy (anonymous):

ty again ^^

OpenStudy (anonymous):

yup